OpenStudy (anonymous):

Use the quadratic formula to solve x2 + 7x + 12 = 0

OpenStudy (anonymous):

The quadratic formula states that the solution x is equal to -b+-(sqrt(b^2-4xaxc))/2a Where a, b, and c correspond to the coefficients in front of the terms in the quadratic. In this case (assuming x2 is x^2), a=1 b=7 c=12 so the values of x are -3 and -4

OpenStudy (campbell_st):

the gereral form of a quadratic is ax^2 +bx + c = 0 in your question a = 1 b = 7 and c = 12 the general quadratic formula is \[x = (-b \pm \sqrt{b^2 - 4ac})/2a\] substitute the values and you will find x = -3 and -4

OpenStudy (anonymous):

x = -3 and x= -4

OpenStudy (anonymous):

ty all for the help